What value of x makes this equation true?
1/5 (2x - 10) + 4x = -3 (1/5x + 4)
Thanks.

Answer by rfer(16322)

2/5x-2+4x=-3/5x-12
2/5x+4x+3/5x=-12+2
5x=-10
x=-10/5
x=-2